OpenStudy (anonymous):

Determine which polynomial is a perfect square trinomial. 49x2 − 28x + 16 9a2 − 30a + 25 25b2 − 45b − 81 16x2 − 24x − 9

OpenStudy (ghazi):

16x2-24x-p= (4x-3)^2

OpenStudy (anonymous):

How did you figure that out so fast? I want to learn too.

OpenStudy (ghazi):

just uses (a-b)^2 and try to fit the above equation in this

OpenStudy (ghazi):

i mean that satisfies a^2-2ab+b^2

jimthompson5910 (jim_thompson5910):

Take the square root of the first and last terms. So for example in \[\Large 49x^2-28x+16\], take the square root of 49x^2 to get 7x and take the square root of16 to get 4 The middle term is either equal to -2*7x*4 = -56x or 2*7x*4 = 56x So either \[\Large 49x^2-56x+16\] or \[\Large 49x^2+56x+16\] are perfect squares This means that \[\Large 49x^2-28x+16\] is NOT a perfect square

OpenStudy (anonymous):

thank you